OpenStudy (anonymous):

16 weeks to seconds dimensional analysis? sort of has to do with math help!!

OpenStudy (anonymous):

\[16\text{ weeks}\times\frac{7\text{ days}}{1\text{ week}}\times\frac{24\text{ hours}}{1\text{ day}}\times\frac{60\text{ minutes}}{1\text{ hour}}\times\frac{60\text{ seconds}}{1\text{ minute}}\]

OpenStudy (anonymous):

\[16\color{red}{\cancel{\text{ weeks}}}\times\frac{7\color{blue}{\cancel{{\text{ days}}}}}{1\color{red}{\cancel{\text{ week}}}}\times\frac{24\color{green}{\cancel{\text{ hours}}}}{1\color{blue}{\cancel{\text{ day}}}}\times\frac{60\cancel{\text{ minutes}}}{1\color{green}{\cancel{\text{ hour}}}}\times\frac{60\text{ seconds}}{1\cancel{\text{ minute}}}\] As you can see, the units are canceled out and we basically have \[16\text{ weeks}=16\times7\times24\times60\times60\text{ seconds}=9,676,800\text{ seconds}\] To check, we can refer to WolframAlpha: http://www.wolframalpha.com/input/?i=convert+16+weeks+to+seconds